OpenStudy (anonymous):

Verify that the function satisfies the three hypotheses of Rolle's Theorem on the given interval. Then find all numbers c that satisfy the conclusion of Rolle's Theorem. f(x) = x^3 − x^2 − 2x + 5, [0, 2]

OpenStudy (anonymous):

Derivative is 3x^2-2x-2 F(2) = 5 F(0) = 5 F(2)=F(0) = 0 3x^2-2x-2=0

OpenStudy (anonymous):

satisfies the hypotheses because any polynomial is continuous and differentiable you last job is to set \[3x^2-2x-2=0\] and solve for \(x\)

OpenStudy (anonymous):

but i am confused, because you wrote \(f(2)=5\) and later wrote \(f(2)=0\) it can't be both